The moon has mass M 7.7 1022 kg and radius r 1.7 106 m. Calculate work done needed to launch a 1500 kg spacecraft from its surface to other surface.


1
Expert's answer
2021-06-29T13:15:59-0400

"W=-\Delta U=-(\frac{GMm}{\infin}-\frac{GMm}{r})=\frac{GMm}{r}=\frac{6.67\cdot10^{-11}\cdot7.7\cdot10^{22}\cdot1500}{1.7\cdot10^6}=4.5\cdot10^9\ (J)" . Answer